OpenStudy (idealist10):

Integrate x/(1+xy) dy?

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Is \(x\) fixed, or are you computing a double integral?

OpenStudy (idealist10):

It's a double integral x/(1+xy) dy dx from 0 to 1 and from 0 to 1.

OpenStudy (anonymous):

\[I=\int_0^1\int_0^1\frac{x}{\color{green}{1+xy}}\,\color{red}{dy}\,dx\] Integrating with respect to \(y\) first, you can fix \(x\) so that you treat it like a constant. Substitute \(\color{green}{u=1+xy}\), then take the differential (remember, \(x\) is fixed), which would be \(\color{blue}{du}=\color{blue}{x}\color{red}{\,dy}\). \[I=\int_0^1\int_1^{1+x}\frac{x}{\color{green}u}\color{blue}{\frac{du}{x}}\,dx\]

OpenStudy (idealist10):

But how to integrate ln abs(1+x) dx?

OpenStudy (anonymous):

With respect to \(x\), you're integrating over the interval \([0,1]\). For \(x>-1\), you have \(\ln|1+x|=\ln(1+x)\), so you have \[I=\int_0^1\ln(1+x)\,dx\]

OpenStudy (idealist10):

Okay, thanks!